Suppose the production function of a firm is given as 1/ 2 1/ 2 X  0.5L K prices of labor and capital is given as $ 5 and $ 10 respectively, and the firm has a constant cost outlay of $ 600.Find the combination of labor and capital that maximizes the firm’s output


1
Expert's answer
2021-09-24T11:01:24-0400


The production function is represented as Q = L"\times"K or

TC = L"\times"K


Where: L represent variable inputs (Labor)

K represents fixed inputs (Capital)

"\frac 12"X = 0.5L K or X = K"\times"L

L = $ 5

K = $ 10


Output maximization is obtained by minimizing cost of production

Therefore:

A combination of minimizing capital and labor is

Marginal Rate of Technical Substitution (MRTS) = "\frac {MP_L}{MP_K}"

Marginal product of labor = "\frac{dx}{dl}" = K

'' '' '' capital =  "\frac {dx}{dk}"= L

Therefore, the marginal rate of technical substitution is = "\frac KL"

but :

X =600 (X) will be represented by Total Cost (TC)

L = $ 5

K = $ 10

"\frac KL"= "\frac{10}{5}"

k= 2L


Substituting in the equation

"\frac 12" TC = 0.5L"\times"K


"\frac12" TC = 0.5L"\times" (2L)


"\frac12" "\times" 600 = "L^2"


300 = "L^2"

L =   "\sqrt 300"

L = 17.32

"\therefore"

300 = 0.5"\times" (17.32)"\times" K

K = 34.64